# Low Complexity Channel Estimation for RIS-Assisted THz Systems with Beam Split

### v1

@@ VI Conclusion
In this paper, we investigated channel estimation in RIS-assisted wideband systems with consideration of beam split effect. We proposed an efficient and low-complexity channel estimation scheme. Specifically, we first proposed a novel CBS-GAMP approach to estimate full CSI within a small subset of SCs. Then, to reduce computational complexity, we proposed a novel BSAD scheme which consists of two direction estimation algorithm, i.e., EnM and DS-MUSIC to acquire angular information at BS and RIS, respectively. With knowledge of angular information, we transformed channel estimation into a simple linear-regression problem and recovered CSI at the remaining SCs using a simple LS algorithm. We demonstrated from numerical results that the proposed scheme can achieve superior NMSE performance and significantly reduce pilot overhead.

### v2

@@ VI Conclusion
In this paper, we investigated channel estimation in RIS-assisted wideband THz systems with consideration of beam split effect. We proposed a two-phase channel estimation scheme to achieve accurate CSI acquisition with reduced computational complexity. Specifically, in Phase I, we introduced a two-step method to achieve full CSI estimation over two SCs. In Phase II, we employed angular information obtained from Phase I to estimate the channels for the remaining SCs using a simple LS algorithm. Theoretical analysis and numerical results demonstrated that our scheme achieved near-oracle performance in terms of NMSE. Moreover, compared to existing solutions, the proposed two-phase strategy significantly reduced computational complexity while maintaining a high level of accuracy in CSI acquisition.